# Optimal sampling of tensor networks targeting wave function’s fast decaying tails

### v1

@@ III Optimal tensor network sampling
IV Benchmarking sampling on applications
IV.1 Quantum state with Gaussian probability distribution with matrix product states
$\mathcal{S}=\frac{t_{\text{standard}}}{t_{\text{OPES}}},$ (21)
with $t_{\text{OPES}}$ and $t_{\text{standard}}$ the computational time for the OPES and standard sampling, respectively. In Fig. 3b, we show that the speedup achieved scales with the threshold $\epsilon$, increasing from one order of magnitude for $\epsilon=10^{-2}$ to three orders of magnitude for $\epsilon=10^{-4}$. An in-depth argument on the scaling and speedup can be found in App. .1, but we have no expected theoretical scaling for a comparison in case of Eq. (19). Finally, the computational time $t$ for OPES sampling is almost independent of the chosen coverage threshold. These results demonstrate the superiority of OPES sampling over traditional sampling when one aims to cover most of the probability distribution.
[…]
IV.2 Quantum image with matrix product states
We proceed and analyze a state obtained after the application of a quantum circuit. In particular, we are not interested in exploring final states of quantum algorithms, since these algorithms are built to be easy to sample to minimize the number of measurements; examples of such algorithms are the Grover or Shor algorithms Grover (1996); Shor (1994). We instead investigate a state obtained "at runtime". To perform operations on images using quantum hardware, we first have to load the image. There are several possible ways of doing so: amplitude encoding, angle encoding, or flexible representation of quantum images (FRQI) Le et al. (2011). We focus on the latter case, encoding images using the algorithm in Ref. Amankwah et al. (2022). The question at hand is if we can reconstruct the image with high accuracy after loading it, using tensor network sampling. This procedure helps in understanding if the image is correctly reproduced even for sizes that are not accessible with the exact state vector simulation. In particular, we choose to encode the quantum tea logo, shown in Fig. 4a. The logo is a $600\times 600$ grey-scale image with a total of $M=360000$ pixels encoded in a $20$-qubit state. We now proceed to a quantitative study of the sampling. In Fig. 4b, we show the computational time $t$ needed to reach a given mean square error MSE with the original image, defined as:
where $\text{P}_{sv}$ is the picture obtained from the complete state vector of the state and $\text{P}_{sm}(t)$ is the sampled picture at time $t$. The index $i$ runs over the pixels of the image. Thus, the MSE is improved by over two orders of magnitude in the figure of merit while saving one order of magnitude in computational time. We also show the logos retrieved from the quantum state at different superiterations of the OPES sampling in Fig. 4a. The FRQI algorithm benefits in our opinion the most from the OPES sampling out of our examples. As we see in the inline of Fig. 4b, the speedup $\mathcal{S}$ of OPES over the standard method is exponential in the MSE, which is agreement with App. .1.
[…]
IV.3 Quantum Ising model with tree tensor networks
As the next example, we consider the quantum Ising model Sachdev (2011) as an example from condensed matter. We use the definition of the quantum Ising model as
[…]
The third example moves from pure states, which we consider throughout all the other examples, to density matrices. We generate finite temperature density matrices for the external fields $g=0.5$ and $g=1.0$ via an imaginary time-dependent variational principle (TDVP) time evolution Haegeman et al. (2016) with a locally purified tensor network (LPTN) Werner et al. (2016), which is then converted into a tree tensor operator (TTO) Reinić (2022). The behavior for the external field resembles the ground state, i.e., OPES shows a greater benefit inside the ferromagnetic phase for the data points with $g=0.5$. Surprisingly, the data indicates a bigger speedup at higher temperatures although the high-temperature limit itself is the completely-mixed state with a flat distribution. In summary, any problem where a lower limit on the unsampled space is required can profit significantly from OPES even in scenarios similar to the condensed matter case across all phases treated here.
IV.4 Crosstalk errors in Rydberg atoms with tree tensor networks
An error analysis is useful for the benchmarking of quantum processing units, e.g., for the state preparation. In such a scenario, we expect to be almost perfectly in the target state, where the remaining probability represents the error. Depending on the fidelity of the preparation, the measurement of a state with error has a small probability. The preparation of the Greenberger-Horne-Zeilinger state (GHZ) with errors including the projective measurement is shown in Ref. Jaschke et al. (2022) and we consider here the aspect of the computational scaling of analyzing such a state. Figure 6a shows the computational time $t$ for an imperfect 64-qubit GHZ state, where errors have been introduced during the protocol due to crosstalk; while the computation time for projective measurements scales with the number of samples for the standard approach, the OPES approach has an almost constant computational time $t$. This almost constant time $t$ is related to the two large probabilities for the states |$\ket{00\ldots 00}$ and |$\ket{11\ldots 11}$. This scenario is one of the most favorable for OPES.
[…]
IV.5 Magic with matrix product states
To stress the generality of the method mentioned in Sec. I, we now generalize the sampling procedure to measure the quantum magic of a quantum state, also known as non-stabilizerness. Quantum magic is an important quantity since it limits the classical simulation of quantum systems for stabilizer simulators Howard et al. (2014); Seddon et al. (2021), similarly to how the entanglement limits the simulations with tensor networks. Quantum magic can be measured through Stabilizer Rényi Entropies (SREs) Leone et al. (2022). Reference Lami and Collura (2023) introduces a technique to measure the SREs through standard sampling of an MPS state. Thus, applying OPES can speed up the process of measuring quantum magic. The magic of a $n$-qubits state using the $m$-order SREs is defined as:
where $\Pi(\vec{\sigma})$ is the probability of measuring a given Pauli string, $\sigma_{i}\in\{\mathbb{I},X,Y,Z\}$. To apply OPES to measure $\mathcal{M}$, we target the $\Pi(\vec{\sigma})$, keeping into account that each measure of probability on the single site $\pi(\sigma_{i})$ branches into four possibilities ($\sigma_{i}\in\{\mathbb{I},X,Y,Z\}$) instead of the two possibilities ($b_{i}=\{\ket{0},\ket{1}\}$) treated up to now. For an extensive explanation of how to measure $\pi(\sigma_{i})$ efficiently with an MPS, please refer to Ref. Lami and Collura (2023). As an example, we compute the order-1 SREs of the state
$-\sin^{2}\phi\log\left(\left|\sin\phi\right|\right).$ (26)
In Fig. 7, we address the computation of $\mathcal{M}_{1}$ for $n=10$ qubits using for OPES $n_{S}=10$ superiterations with $n_{r}=100$ samples each. The results are mediated over $10$ realization of the computations. We use for the standard method the same number of samples. We notice that, due to the flatness of the distribution, OPES gives no relevant improvement when $\phi=\frac{\pi}{64}\sim 0$. This result is to be expected, as already discussed for the paramagnetic phase of the quantum Ising model in Sec. IV.3. However, we can observe both an improvement in the estimation of the magic and a speedup in the other two cases, i.e. $\phi=\frac{\pi}{8}$ and $\phi=\frac{\pi}{4}$. Optimal tensor network sampling is twice as fast as the standard method, and the estimation of the quantum magic converges much faster with respect to the standard method.
V Conclusions
Overall, we have introduced the optimal tensor network sampling (OPES), a sampling method for tensor network states that avoids sampling known intervals and caches information for optimal performance. From the computational statistics point of view, the algorithm represents an extension to the inverse transform sampling, which is used to sample pseudo-random numbers from an arbitrary probability distribution function by knowing its inverse probability distribution function. The extension consists in the fact that we never fully know the cumulative distribution function or its inverse of the quantum state represented as a tensor network. The method is applicable to all isometric tensor network states as well as exact diagonalization. We have investigated several quantum states and compared the performance of the standard and OPES methods in characterizing them; these results are obtained for matrix product states (MPS), tree tensor networks (TTN), and tree tensor operators (TTO).
In detail, we have shown that OPES offers a significant speedup compared to the standard method, especially when covering most of the probability distribution is demanded. On the one hand, this speedup is due to the efficient caching of previous information employing the MPS or TTN; on the other hand, we work with the underlying probability intervals of the full Hilbert space which allow us to exclude previously explored intervals from future superiterations. For example, OPES leads to a speedup of more than an order of magnitude while achieving lower errors versus the standard sampling for applications as the flexible representation of quantum images (FRQI) algorithm. We have shown how the approach is generalizable to the measure of magic with matrix product state, and its speedup over standard methods. The speedup in computation time also leads to a reduced integrated energy consumption of the CPU as our additional data in Ballarin et al. (2024b) shows for one example.
Future directions of this approach for sampling can go into the direction of comparing parallelized versions, optimizing the caching strategy, or exploring the potential benefit in other algorithms and applications. While the standard sampling trivially parallelizes via data parallelism, OPES can, for example, parallelize over the groups of the sorted random numbers. For the application side, there exists a plethora of candidates with potential speedup where the examples given within this manuscript give a first intuition which applications benefit most from OPES, i.e., quantum states with a decaying probability distribution. Furthermore, since OPES only needs the computation of expectation values of local operators and marginal probabilities it could be applied to any isometric tensor network such as isoPEPS Zaletel and Pollmann (2020). Similarly, the strings of projective measurements can be naturally replaced with Positive Operator-Valued Measures (POVM).
Code and data availability -- The code implementing the algorithm for both MPS and TTNs is distributed through the quantum tea leaves Ballarin et al. (2024a) python package of Quantum TEA, while the code and data to generate the plots of this work is available via zenodo Ballarin et al. (2024c, d). All figures are available together with metadata under a CC-BY license at Ballarin et al. (2024b).
Acknowledgments -- We thank Nora Reinić for providing finite-temperature states in the tree tensor operator representation. We thank Giovanni Cataldi, Marcello Dalmonte, Martina Frau, Marco Rigobello, Ilaria Siloi, Poetri Tarabunga, Marco Tesoro, and Emanuele Tirrito for discussions and feedback. We acknowledge financial support from the Italian Ministry of University and Research (MUR) via PRIN2022 project TANQU, and the Departments of Excellence grant 2023-2027 Quantum Frontiers; from the European Union via H2020 projects EuRyQa, and TEXTAROSSA, the QuantERA projects QuantHEP and T-NISQ, and the Quantum Flagship project Pasquans2, from the German Federal Ministry of Education and Research (BMBF) via the funding program quantum technologies -- from basic research to market -- project QRydDemo, and from the World Class Research Infrastructure -- Quantum Computing and Simulation Center (QCSC) of Padova University. This work was performed in part at the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452; the participation of D. J. at the Aspen Center for Physics was supported by the Simons Foundation. We also acknowledge computation time supported support by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant no INST 40/575-1 FUGG (JUSTUS 2 cluster), the Atos Dibona cluster via TEXTAROSSA, as well as computation time on Cineca’s Galileo100 machine.
Supplemental material
The appendices focus on providing additional information which is necessary to implement the OPES algorithm efficiently. Section .1 gives a brief example on the scaling of both algorithms. Section .2 extends the pedagogical, simplified algorithm to measuring multiple shots in superiterations, while Sec. .3 explains the tensor network aspects and how to cache information for the tensor network.
.1 Scaling of the algorithms
We layout a brief calculation of how one can estimate the scaling of the two algorithms and focus on the scaling for the coverage $\mathcal{C}$. Recall that the threshold is $\epsilon=1-\mathcal{C}$. During the derivations, we have to make assumptions on the probability distribution. As a reference distribution, we introduce the series of probabilities
@@ V Conclusions
.2 Implementation of the algorithm
@@
Section III contains already one version of the algorithm; now, we extend the description to cover optional features like multiple shots, superiterations, or cache access. The algorithm takes as input the number of random numbers to be sampled $n_{r}$, the tensor network, and the set $\Delta=[0,1]\setminus\overline{\Delta}=\{\delta_{1},\ldots,\delta_{m}\}$ of $m$ known intervals $\delta_{\alpha}$. A superiteration of the algorithm has the following steps:
@@
    Repeat step 5 until all the sites of the system are measured. Further details for this step are explained in Sec. III and remain valid here.
@@
    Remove the measured probability interval from our sampling interval, i.e. $\Delta=\Delta\setminus\delta_{\alpha}$. Each new sampled number $u_{j}~{}\sim U(\Delta)$ explores a different probability interval $\delta_{\beta\neq\alpha}$. Update the cache with the corresponding information and tensors.
@@
.3 Details of the tensor networks implementation

### v2

@@ 3 Optimal tensor network sampling
4 Theoretical scalings
We proceed now to analyze the optimality of OPES and its theoretical scaling in treatable scenarios, i.e., for a flat and a decaying distribution.
4.1 Uniform distribution
In a uniform distribution, the probability of each state is $p=1/D$ where $D$ is the dimension of the Hilbert space. Say that we sample anyway and search for some state |$\ket{\psi}$, we know that the probability to have sampled |$\ket{\psi}$ after $n$ iterations is
$p(n)$ $=1-\left(1-\frac{1}{D}\right)^{n}=\frac{D^{n}-(D-1)^{n}}{D^{n}}\,.$ (19)
Instead with OPES, we have
$p_{OPES}(n)$ $=1-\prod_{i=0}^{n-1}\left(1-\frac{1}{D-i}\right)=\frac{n}{D}\,,$ (20)
which reflects that after $n=D$ we reach the state in any scenario. We rewrite it as ratio
$\frac{p}{p_{OPES}}$ $=\frac{1}{n}\sum_{i=0}^{n-1}\frac{(-1)^{i}\begin{pmatrix}n\\ n-1-i\end{pmatrix}}{D^{i}}\,.$ (21)
The ratio does not tell us much and is in the interval $[p(D),1]$. We might have to change the question as ask how many samples are required to find the state |$\ket{\psi}$ with a probability $q$. Then, we have
$n(p(n)\geq q)$ $=\frac{\log(1-q)}{\log\left(1-\frac{1}{D}\right)}\,,$ (22)
$n_{OPES}(p_{OPES}(n)\geq q)$ $=q\cdot D\,.$ (23)
The expression for standard sampling diverges for $q\to 1$. The question now is how fast the function diverges. The divergence logarithmically with $q$.
Figure 3: Establishing scaling for uniform distribution. Speedup for OPES sampling all space up to $\epsilon$ for a uniform distribution and 50 qubits. We observe a log speedup with $\epsilon$.
To get an equal approach of the unsampled space $\epsilon$, we can analyze the Poisson distribution $P_{\lambda=n/D}(k=0)=\lambda^{k}/k!\cdot e^{-\lambda}$ to get the number of states which have not been sampled, i.e., $D\cdot e^{-n/D}$:
$n(\epsilon)$ $=-D\cdot\log(\epsilon)\,,$ (24)
$n_{OPES}(\epsilon)$ $=D(1-\epsilon)\,.$ (25)
Assuming the overhead for bookkeeping is minimal, we have an OPES speedup of
$S(\epsilon)$ $=\frac{-\log(\epsilon)}{(1-\epsilon)}\,.$ (26)
The log-speedup of the uniform distribution can be seen as lower bound for the speedop with OPES as it is the most difficult one to sample up to an $\epsilon$.
4.2 Exponentially decaying distribution
We define the exponentially decaying distribution as follows:
@@ 4.2 Exponentially decaying distribution
5 Benchmarking sampling on applications
5.1 Quantum state with Gaussian probability distribution with matrix product states
$\mathcal{S}=\frac{t_{\text{standard}}}{t_{\text{OPES}}},$ (34)
with $t_{\text{OPES}}$ and $t_{\text{standard}}$ the computational time for the OPES and standard sampling, respectively. In Fig. 5b, we show that the speedup achieved scales with the threshold $\epsilon$, increasing from one order of magnitude for $\epsilon=10^{-2}$ to three orders of magnitude for $\epsilon=10^{-4}$. An in-depth argument on the scaling and speedup can be found in Sec. 4, but we have no expected theoretical scaling for a comparison in case of Eq. (32). Finally, the computational time $t$ for OPES sampling is almost independent of the chosen coverage threshold. These results demonstrate the superiority of OPES sampling over traditional sampling when one aims to cover most of the probability distribution.
[…]
5.2 Quantum image with matrix product states
We proceed and analyze a state obtained after the application of a quantum circuit. In particular, we are not interested in exploring final states of quantum algorithms, since these algorithms are built to be easy to sample to minimize the number of measurements; examples of such algorithms are the Grover or Shor algorithms [69, 70]. We instead investigate a state obtained "at runtime". To perform operations on images using quantum hardware, we first have to load the image. There are several possible ways of doing so: amplitude encoding, angle encoding, or flexible representation of quantum images (FRQI) [71]. We focus on the latter case, encoding images using the algorithm in Ref. [72]. The question at hand is if we can reconstruct the image with high accuracy after loading it, using tensor network sampling. This procedure helps in understanding if the image is correctly reproduced even for sizes that are not accessible with the exact state vector simulation. In particular, we choose to encode the quantum tea logo, shown in Fig. 6a. The logo is a $600\times 600$ grey-scale image with a total of $M=360000$ pixels encoded in a $20$-qubit state. We now proceed to a quantitative study of the sampling. In Fig. 6b, we show the computational time $t$ needed to reach a given mean square error MSE with the original image, defined as:
where $\text{P}_{sv}$ is the picture obtained from the complete state vector of the state and $\text{P}_{sm}(t)$ is the sampled picture at time $t$. The index $i$ runs over the pixels of the image. Thus, the MSE is improved by over two orders of magnitude in the figure of merit while saving one order of magnitude in computational time. We also show the logos retrieved from the quantum state at different superiterations of the OPES sampling in Fig. 6a. The FRQI algorithm benefits in our opinion the most from the OPES sampling out of our examples. As we see in the inline of Fig. 6b, the speedup $\mathcal{S}$ of OPES over the standard method is exponential in the MSE, which is agreement with Sec. 4.
[…]
5.3 Quantum Ising model with tree tensor networks
As the next example, we consider the quantum Ising model [73] as an example from condensed matter. We use the definition of the quantum Ising model as
[…]
The third example moves from pure states, which we consider throughout all the other examples, to density matrices. We generate finite temperature density matrices for the external fields $g=0.5$ and $g=1.0$ via an imaginary time-dependent variational principle (TDVP) time evolution [74] with a locally purified tensor network (LPTN) [45], which is then converted into a tree tensor operator (TTO) [49]. The behavior for the external field resembles the ground state, i.e., OPES shows a greater benefit inside the ferromagnetic phase for the data points with $g=0.5$. Surprisingly, the data indicates a bigger speedup at higher temperatures although the high-temperature limit itself is the completely-mixed state with a flat distribution. In summary, any problem where a lower limit on the unsampled space is required can profit significantly from OPES even in scenarios similar to the condensed matter case across all phases treated here.
5.4 Crosstalk errors in Rydberg atoms with tree tensor networks
An error analysis is useful for the benchmarking of quantum processing units, e.g., for the state preparation. In such a scenario, we expect to be almost perfectly in the target state, where the remaining probability represents the error. Depending on the fidelity of the preparation, the measurement of a state with error has a small probability. The preparation of the Greenberger-Horne-Zeilinger state (GHZ) with errors including the projective measurement is shown in Ref. [14] and we consider here the aspect of the computational scaling of analyzing such a state. Figure 8a shows the computational time $t$ for an imperfect 64-qubit GHZ state, where errors have been introduced during the protocol due to crosstalk; while the computation time for projective measurements scales with the number of samples for the standard approach, the OPES approach has an almost constant computational time $t$. This almost constant time $t$ is related to the two large probabilities for the states |$\ket{00\ldots 00}$ and |$\ket{11\ldots 11}$. This scenario is one of the most favorable for OPES.
[…]
5.5 Magic with matrix product states
To stress the generality of the method mentioned in Sec. 1, we now generalize the sampling procedure to measure the quantum magic of a quantum state, also known as non-stabilizerness. Quantum magic is an important quantity since it limits the classical simulation of quantum systems for stabilizer simulators [75, 76], similarly to how the entanglement limits the simulations with tensor networks. Quantum magic can be measured through Stabilizer Rényi Entropies (SREs) [77, 78]. Reference [56] introduces a technique to measure the SREs through standard sampling of an MPS state. Thus, applying OPES can speed up the process of measuring quantum magic. The magic of a $n$-qubits state using the $m$-order SREs is defined as:
where $\Pi(\vec{\sigma})$ is the probability of measuring a given Pauli string, $\sigma_{i}\in\{\mathbb{I},X,Y,Z\}$. To apply OPES to measure $\mathcal{M}$, we target the $\Pi(\vec{\sigma})$, keeping into account that each measure of probability on the single site $\pi(\sigma_{i})$ branches into four possibilities ($\sigma_{i}\in\{\mathbb{I},X,Y,Z\}$) instead of the two possibilities ($b_{i}=\{\ket{0},\ket{1}\}$) treated up to now. For an extensive explanation of how to measure $\pi(\sigma_{i})$ efficiently with an MPS, please refer to Ref. [56]. As an example, we compute the order-1 SREs of the state
$-\sin^{2}\phi\log\left(\left|\sin\phi\right|\right).$ (39)
In Fig. 9, we address the computation of $\mathcal{M}_{1}$ for $n=10$ qubits using for OPES $n_{S}=10$ superiterations with $n_{r}=100$ samples each. The results are mediated over $10$ realization of the computations. We use for the standard method the same number of samples. We notice that, due to the flatness of the distribution, OPES gives no relevant improvement when $\phi=\frac{\pi}{64}\sim 0$. This result is to be expected, as already discussed for the paramagnetic phase of the quantum Ising model in Sec. 5.3. However, we can observe both an improvement in the estimation of the magic and a speedup in the other two cases, i.e., $\phi=\frac{\pi}{8}$ and $\phi=\frac{\pi}{4}$. Optimal tensor network sampling is twice as fast as the standard method, and the estimation of the quantum magic converges much faster with respect to the standard method.
6 Conclusions
Overall, we have introduced the optimal tensor network sampling (OPES), a sampling method for tensor network states that avoids sampling known intervals and caches information for optimal performance. From the computational statistics point of view, the algorithm represents an extension to the inverse transform sampling, which is used to sample pseudo-random numbers from an arbitrary probability distribution function by knowing its inverse probability distribution function. The extension consists in the fact that we never fully know the cumulative distribution function or its inverse of the quantum state represented as a tensor network. The method is applicable to all isometric tensor network states as well as exact diagonalization. We have investigated several quantum states and compared the performance of the standard and OPES methods in characterizing them; these results are obtained for matrix product states (MPS), tree tensor networks (TTN), and tree tensor operators (TTO). We compare to the standard sampling method. An enhanced standard sampling, as available in Ref. [79], can share some of the features of OPES as the speedup at the cost of sampling as well as storing more random numbers and loosing convenience in the numerical implementation.
In detail, we have shown that OPES offers a significant speedup compared to the standard method, especially when covering most of the probability distribution is demanded. On the one hand, this speedup is due to the efficient caching of previous information employing the MPS or TTN; on the other hand, we work with the underlying probability intervals of the full Hilbert space which allow us to exclude previously explored intervals from future superiterations. For example, OPES leads to a speedup of more than an order of magnitude while achieving lower errors versus the standard sampling for applications as the flexible representation of quantum images (FRQI) algorithm. We have shown how the approach is generalizable to the measure of magic with matrix product state, and its speedup over standard methods. The speedup in computation time also leads to a reduced integrated energy consumption of the CPU as our additional data in Ref. [80] shows for one example.
Future directions of this approach for sampling can go into the direction of comparing parallelized versions, optimizing the caching strategy, or exploring the potential benefit in other algorithms and applications. While the standard sampling trivially parallelizes via data parallelism, OPES can, for example, parallelize over the groups of the sorted random numbers. For the application side, there exists a plethora of candidates with potential speedup where the examples given within this manuscript give a first intuition which applications benefit most from OPES, i.e., quantum states with a decaying probability distribution. Furthermore, since OPES only needs the computation of expectation values of local operators and marginal probabilities it could be applied to any isometric tensor network such as isoPEPS [61]. Similarly, the strings of projective measurements can be naturally replaced with Positive Operator-Valued Measures (POVM).
Code and data availability -- The code implementing the algorithm for both MPS and TTNs is distributed through the quantum tea leaves [65] python package of Quantum TEA, while the code and data to generate the plots of this work is available via zenodo [81, 82]. All figures are available together with metadata under a CC-BY license at [80].
Acknowledgments -- We thank Nora Reinić for providing finite-temperature states in the tree tensor operator representation. We thank Giovanni Cataldi, Marcello Dalmonte, Martina Frau, Marco Rigobello, Ilaria Siloi, Poetri Tarabunga, Marco Tesoro, and Emanuele Tirrito for discussions and feedback. We acknowledge financial support from the Italian Ministry of University and Research (MUR) via PRIN2022 project TANQU, and the Departments of Excellence grant 2023-2027 Quantum Frontiers; from the European Union via H2020 projects EuRyQa, and TEXTAROSSA, the QuantERA projects QuantHEP and T-NISQ, and the Quantum Flagship project Pasquans2, INFN project Quantum, from the German Federal Ministry of Education and Research (BMBF) via the funding program quantum technologies -- from basic research to market -- project QRydDemo, and from the World Class Research Infrastructure -- Quantum Computing and Simulation Center (QCSC) of Padova University. This work was performed in part at the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452; the participation of D. J. at the Aspen Center for Physics was supported by the Simons Foundation. We also acknowledge computation time supported support by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant no INST 40/575-1 FUGG (JUSTUS 2 cluster), the Atos Dibona cluster via TEXTAROSSA, as well as computation time on Cineca’s Galileo100 machine.
Supplemental material
The appendices focus on providing additional information which is necessary to implement the OPES algorithm efficiently. Section A extends the pedagogical, simplified algorithm to measuring multiple shots in superiterations, while Sec. B explains the tensor network aspects and how to cache information for the tensor network.
Appendix A Implementation of the algorithm
@@
Section 3 contains already one version of the algorithm; now, we extend the description to cover optional features like multiple shots, superiterations, or cache access. The algorithm takes as input the number of random numbers to be sampled $n_{r}$, the tensor network, and the set $\Delta=[0,1]\setminus\overline{\Delta}=\{\delta_{1},\ldots,\delta_{m}\}$ of $m$ known intervals $\delta_{\alpha}$. A superiteration of the algorithm has the following steps:
@@
    Repeat step 5 until all the sites of the system are measured. Further details for this step are explained in Sec. 3 and remain valid here.
@@
    Remove the measured probability interval from our sampling interval, i.e., $\Delta=\Delta\setminus\delta_{\alpha}$. Each new sampled number $u_{j}~{}\sim U(\Delta)$ explores a different probability interval $\delta_{\beta\neq\alpha}$. Update the cache with the corresponding information and tensors.
@@
A.1 Uniformly sampling from $\Delta$
In this section, we discuss how the $n_{r}$ uniform random variables $\{u_{j}\}_{j=1,\dots,n_{r}}$ from the interval $\Delta$ are sampled. First, we define $\Delta_{i}$ the $M$ continuous intervals componing $\Delta$, i.e. $\Delta=\bigcup_{i=1}^{M}\Delta_{i}$. We define the probability related to the interval $\Delta_{i}$ as $p_{i}$. The technique to sample $u_{j}$ is as follows:
  1. 1.
    Determine how many samples $n_{r}^{i}$ are sampled from each interval $\Delta_{i}$, with $\sum_{i=1}^{M}n_{r}^{i}=n_{r}$. This task can be easily done by sampling $n_{r}$ numbers uniformly in $U(0,1)$, and identify the sample $u_{k}\sim U(0,1)$ as belonging to $\Delta_{i}$ if:
    $\widetilde{p}_{i-1}\leq u_{k}<\widetilde{p}_{i},$ (49)
    with $\widetilde{p}_{0}=0$ and $\widetilde{p}_{i}=p_{i}/\sum_{i=1}^{M}p_{i}$.
  2. 2.
    Sample $n_{r}^{i}$ random numbers from the interval $\Delta_{i}$:
    $\{u_{j}\}_{j=1,\dots,n_{r}}$ $=\bigcup_{i=1}^{M}\;\big{\{}u_{j^{\prime}}\sim U(\Delta_{i})\big{\}}_{j^{\prime}=1,\dots,n_{r}^{i}}$ (50)
    $=\bigcup_{j=1}^{n_{r}}\;\big{\{}u_{j}\sim U(\Delta)\big{\}}$ (51)
The resulting $\{u_{j}\}_{j=1,\dots,n_{r}}$ are uniformly distributed over $\Delta$. We stress that the computational complexity $C_{\Delta}$ of sampling $n_{r}$ random numbers from a discontinuous interval $\Delta$ with $M$ continuous intervals $\Delta_{i}$ scales as:
$C_{\Delta}=O\big{(}n_{r}M\big{)},$ (52)
since we just need to generate $2n_{r}$ random numbers, but we have to check up to $M$ conditions on $n_{r}$ of them.